OpenStudy (anonymous):

Greta had 5/8 pound of ham. She used 1/3 pound for sandwiches. How many pounds of ham does she have left? A.5/24 lb B.7/24 lb C.1/3 lb D.4/5 lb

jigglypuff314 (jigglypuff314):

5/8 - 1/3 = and remember that in order to add/subtract you have to have same denominators

jigglypuff314 (jigglypuff314):

5/8 = 15/24 1/3 = 8/24 then you can subtract 15/24 - 8/24 =

jigglypuff314 (jigglypuff314):

-.- no, it would Not be D

jigglypuff314 (jigglypuff314):

remember that in order to add/subtract you have to have Same denominators otherwise it would be like saying 1/2 of a pizza, I ate 1/4. now I have -1/2 ?!

OpenStudy (anonymous):

Yup :3

jigglypuff314 (jigglypuff314):

good :)